OpenStudy (anonymous):

Find an equation of the line tangent to teh curve at the given point. y=x^2 * e^-x , (1, 1/e)

OpenStudy (anonymous):

take the derivative using the product rule

OpenStudy (anonymous):

you get \[y'=2xe^{-x}-x^2e^{-x}\] and then replace x by 1 to get the slope

OpenStudy (anonymous):

\[f'(1)=\frac{2}{e}-\frac{1}{e}=\frac{1}{e}\] then use point slope formula

OpenStudy (anonymous):

im confused about how u got 2/e

OpenStudy (kira_yamato):

Just plug in x = 1 into the f'(x) or y'

OpenStudy (anonymous):

i got \[\frac{2}{e}\] because the first term of the derivative is \[2xe^{-x}\] and if i replace x by 1 i get \[2\times 1\times e^{-1}=2e^{-1}=\frac{2}{e}\]

OpenStudy (anonymous):

so is that the slope (2/e) that i use 2 plug into teh slope-interept equation? y - (1/e) = 2/e (x- 1)

OpenStudy (anonymous):

y = 2/e(x) - 1/e

OpenStudy (anonymous):

no the slope is not \[\frac{2}{e}\] it is \[\frac{1}{e}\]

OpenStudy (anonymous):

i thought wen we evaluate teh derivative at the x value that wud b the slope...how is 1/e the slope

OpenStudy (anonymous):

plz explain..i know im slow at math